### v1

Abstract
Precise control of signal propagation in modular neural networks represents a fundamental challenge in computational neuroscience. We establish a framework for identifying optimal control nodes that maximize stimulus transmission between weakly coupled neural populations. Using spiking stochastic block model networks, we systematically compare driver node selection strategies — including random sampling and topology-based centrality measures (degree, betweenness, closeness, eigenvector, harmonic, and percolation centrality) — to determine minimal control inputs for achieving inter-population synchronization.
Targeted stimulation of just 10–20% of the most central neurons in the source population significantly enhances spiking propagation fidelity compared to random selection ($p<0.001$). This approach yields a $\sim$2.7-fold increase in signal transfer efficiency at critical inter-module connection densities ($p_{\mathrm{inter}}=0.04-0.07$). These findings establish a theoretical foundation for precision neuromodulation in biological neural systems and neurotechnology applications.
Keywords: network control · neural synchronization · centrality measures · neuromodulation · stochastic block model
1 Introduction
Neural synchronization between distinct brain regions underlies critical cognitive processes including perception, attention, and memory formation [4],[10]. Disruptions in inter-regional synchronization are implicated in neurological disorders such as epilepsy, Alzheimer’s disease, and schizophrenia [9]. Despite its significance, the structural determinants of synchronization propagation, particularly the role of nodal centrality and inter-areal connectivity, remain incompletely characterized.
Identification of minimal driver inputs required to achieve synchronization offers dual benefits: advancing fundamental understanding of brain network principles and developing targeted therapies for synchronization pathologies. We address this challenge through stochastic block model (SBM) networks, building upon prior work in neural network control [2]. Our approach evaluates multiple centrality metrics to determine topological predictors of effective driver nodes.
This research extends beyond previous studies in three key aspects: (1) systematic comparison of six centrality measures for control node identification, (2) explicit quantification of propagation efficiency in spiking neural networks, and (3) analysis of inter-cluster connectivity effects on synchronization transfer. We hypothesize that driver node selection based on specific centrality metrics and inter-cluster connectivity significantly enhances synchronization efficiency in SBM neural networks.
2 Methods
2.1 Network Architecture and Topology Modification
Neural populations were modeled as undirected stochastic block model (SBM) networks [7] with two equally sized clusters ($N=500$ neurons, $|V_{t}|=|V_{o}|=250$, where $|V_{t}|$ and $|V_{o}|$ are the sizes of the stimulated and measured populations of neurons). The adjacency matrix $A$ was constructed with intra-cluster connection probability $p_{\mathrm{intra}}=0.15$ and inter-cluster probability $p_{\mathrm{inter}}$ systematically varied from 0.01 to 0.1. The signal was applied to neurons within the top 10%, 15%, or 20% centrality percentiles ($p_{\mathrm{input}}$) or to 10–20% of randomly selected neurons within the cluster $V_{t}$ (Proxy stimulation strategy and Random baseline). To overcome SBM’s inherent homogeneity and in order to introduce topological diversity, we implemented a centrality boosting procedure defined by:
$\Delta k(u)=\left\lceil(\beta-1)\frac{c(u)}{\max_{v\in S}c(v)}k_{\mathrm{ext}}(u)\right\rceil,$ (1)
where:
  • $c(u)$ is centrality value of the $u$ vertex
  • $\beta=1.5$ (50% centrality boost)
  • $k_{\mathrm{ext}}(u)$ is the existing inter-cluster degree
  • $S$ is the set of selected driver neurons
Network density was preserved via compensatory edge removal using the constraint:
$\sum_{u\in V_{t}}k_{\mathrm{ext}}(u)=\lfloor p_{\mathrm{inter}}\cdot|V_{t}|\cdot|V_{o}|\rfloor.$ (2)
Four distinct control node selection strategies were implemented:
  1. I.
    Top Centrality: Select $n_{\mathrm{boost}}$ neurons in $V_{t}$ with maximal $c(u)$
  2. II.
    Proxy Stimulation: Random selection of neurons in $V_{t}$ from neighbors of top central nodes in $V_{o}$
  3. III.
    Random Baseline: Uniform random selection from $V_{t}$
2.2 Centrality Measure Computation
We evaluated six topological centrality measures that quantify distinct aspects of node influence in networks. For graph $G=(V,E)$ with adjacency matrix $A$ ($A_{uv}=1$ if edge exists, 0 otherwise), $n=|V|$ nodes, and shortest path distance $d(u,v)$:
Degree Centrality
Quantifies direct connectivity influence:
$C_{d}(u)=\sum_{v\in V}A_{uv}.$ (3)
It represents the number of immediate neighbors, indicating local influence capacity.
Betweenness Centrality
Measures brokerage potential in information flow:
$C_{b}(u)=\sum_{\begin{subarray}{c}s\neq u\neq t\\ s,t\in V\end{subarray}}\frac{\sigma_{st}(u)}{\sigma_{st}},$ (4)
where $\sigma_{st}$ is the total shortest paths between $s$ and $t$, and $\sigma_{st}(u)$ is the paths traversing $u$. This centrality identifies nodes functioning as critical bridges.
Closeness Centrality
Characterizes global integration efficiency:
[…]
Nodes with minimal average path distance to others can rapidly disseminate information.
Eigenvector Centrality
Quantifies influence through connectedness to influential neighbors:
$\mathbf{x}=\lambda_{\max}^{-1}A\mathbf{x}.$ (6)
This equation is solved via power iteration to convergence ($\Delta x<10^{-6}$), where $\lambda_{\max}$ is the largest eigenvalue of $A$. This captures recursive importance propagation.
Percolation Centrality
Models dynamic spreading importance:
$C_{p}(u)=\frac{1}{n-2}\sum_{\begin{subarray}{c}s\neq u\neq r\\ s,r\in V\end{subarray}}\frac{\sigma_{sr}(u)}{\sigma_{sr}}\frac{x_{s}^{(t)}}{\sum_{i\in V}x_{i}^{(t)}-x_{u}^{(t)}},$ (7)
where $x_{i}^{(t)}$ is the activation probability at time $t$ (simulated over $t=1$ to $10$). This extends betweenness centrality to time-varying processes.
Harmonic Centrality
Robust closeness variant for disconnected graphs:
$C_{h}(u)=\frac{1}{n-1}\sum_{v\neq u}\frac{1}{d(u,v)}\quad\left(\frac{1}{d(u,v)}:=0\text{ if }d(u,v)=\infty\right).$ (8)
This centrality penalizes unreachable nodes less severely than standard closeness.
All measures were min-max normalized to $[0,1]$ before analysis. Computations used NetworkX 3.0 [6] with Brandes’ algorithm for path-based metrics [3].
2.3 Neural Dynamics Model
We consider spiking neural network consisting of leaky integrate-and-fire (LIF) neurons [1]. As was mentioned in the previous section, the whole network was divided into two equal populations. The connections between individual neurons within and between populations were established according to the desired adjacency matrix.
Each of the two populations contained $80\%$ excitatory and $20\%$ inhibitory neurons. The equations for the membrane potential of a single neuron were written as follows:
$\tau\frac{dv}{dt}=v_{\mathrm{rest}}-v+R\,I_{\mathrm{ext}}+v_{n}+v_{e}+v_{i},$ (9)
where $v_{\mathrm{rest}}$ is the steady-state value of the membrane potential, $R$ is the equivalent membrane resistance, $\tau$ is the characteristic time constant, and $v_{e}$ and $v_{i}$ are changes in the membrane potential caused by synaptic currents from contacting excitatory ($e$) and inhibitory ($i$) neurons:
$v_{e}$ $\mathrel{+}=\sum_{e}w_{e}\,\delta\bigl{(}t-t_{e}\bigr{)},$ (10)
$v_{i}$$\mathrel{-}=\sum_{i}w_{i}\,\delta\bigl{(}t-t_{i}\bigr{)}.$ (11)
To simulate background neural activity, all neurons received independent stimuli $v_{n}$ with a Poisson distribution. Each spike from the Poisson generator causes an instantaneous increase in membrane potential, which can be represented as the sum of delta functions:
$v_{n}\mathrel{+}=\sum_{n}w_{n}\,\delta\bigl{(}t-t_{n}\bigr{)},$ (12)
where $t_{n}$ is the time of arrival of the spikes. We assumed that
$w_{e}=w_{i}=w_{n}=w.$
In addition to noise from external influences, the driver neurons received an additional periodic signal:
$I_{\mathrm{ext}}=I_{0}\,\sin\bigl{(}2\pi ft+\phi\bigr{)}.$ (13)
In our setting, driver neurons were chosen only amongst the excitatory neurons of population $V_{t}$.
The neuron generated a spike when $v(t)$ reached the excitation threshold $v_{\mathrm{th}}$, after which the potential was instantly reset to $v_{\mathrm{reset}}$. Formally, this is described by the condition
$v(t)\geq v_{\mathrm{th}}\quad\longrightarrow\quad v(t)=v_{\mathrm{reset}}.$
All the parameter values used in the simulations are shown in Table 1.
Table 1: Neuron dynamics simulation parameters
Parameter Symbol Value
[…]
2.4 Numerical Experiments and Analysis
Simulation Protocol
  1. 1.
    Network generation (SBM with specified $p_{\mathrm{intra}}$, $p_{\mathrm{inter}}$, 20 times for each parameters pair)
  2. 2.
    Centrality computation and node selection
  3. 3.
    Topology modification ($\Delta k$ edges added)
  4. 4.
    Density compensation (random edge removal)
  5. 5.
[…]
  6. 6.
    100 ms warm-up period excluded from analysis
3 Results
We ran several experiments, in which we choose specific driver nodes in population $V_{t}$ and measured neural activity in population $V_{o}$. We tested several experimental set-ups:
  • network with and without boosting;
  • selecting driver neurons within the population $V_{t}$ among the nodes with maximum centrality measures;
  • random driver neurons of cluster $V_{t}$ that are also the neighbors of the most central neurons of the cluster $V_{o}$;
  • random driver neurons of cluster $V_{t}$ or random baseline.
These topological control strategies demonstrated different efficiencies. In the absence of centrality boosting, driver node selection strategy significantly impacted the activity propagation efficiency. Crucially, stimulation of top-centrality nodes substantially outperformed proxy stimulation across all parameter set elements (see Fig. 2).
Figure 1: The distribution of centrality metrics values for network nodes is shown before boosting (Original) and after boosting (Modified). In the boosted case, a distinct tail emerges representing the nodes with significantly increased centrality of the selected type
Figure 2: Propagation efficiency in non-boosted networks. Performance surfaces show significant divergence between top-centrality (colored solid) and proxy stimulation strategies (colored shaded) and random baseline (pink solid).
Network Control Dynamics in Structural Connectomes: Non-boosted Network Performance
Our analysis reveals fundamental constraints in stimulus propagation through structural connectomes. Below the critical percolation threshold ($p_{\mathrm{inter}}<0.04$), no strategy elicited significant propagation – indicating a fundamental physical constraint on information transmission in sparse connectomes. Beyond this threshold, strategy-dependent effects emerged:
  1. 1.
    Medium connectivity regime ($p_{\mathrm{inter}}=0.07$): Stimulation of top eigenvector centrality nodes achieved moderate propagation ($0.64\pm 0.24\leavevmode\nobreak\ \text{Hz}$), consistent with their role in global network integration. Strikingly, proxy stimulation targeting betweenness-central nodes near hub regions showed severely attenuated efficacy ($0.15\pm 0.14\leavevmode\nobreak\ \text{Hz}$), representing a 76% reduction versus optimal targets. This performance approached the random baseline ($0.01\pm 0.01\leavevmode\nobreak\ \text{Hz}$), demonstrating that anatomical adjacency to hubs is insufficient for effective control.
  2. 2.
    High connectivity regime ($p_{\mathrm{inter}}=0.10$): Degree-centrality targeting maximized propagation ($3.33\pm 0.98\leavevmode\nobreak\ \text{Hz}$), leveraging local integration capacity. Proxy stimulation using degree-central nodes near hubs remained markedly inefficient ($0.57\pm 0.51\leavevmode\nobreak\ \text{Hz}$, 82% reduction). The persistent strategy gap across densities suggests that conventional hub-centric approaches neglect critical control mechanisms beyond local connectivity.
Topological Enhancement Reshapes Control Landscapes Centrality boosting amplified global propagation efficiency by 2.7$\times$ ($p<10^{-10}$) and fundamentally altered network dynamics:
  1. 1.
    Threshold modulation: Boosted networks exhibited significant propagation at $p_{\mathrm{inter}}=0.07$below the non-boosted percolation threshold. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
  2. 2.
    Convergence of control strategies:
    • At $p_{\mathrm{inter}}=0.07$, closeness-centrality targeting dominated ($2.06\pm 0.44\leavevmode\nobreak\ \text{Hz}$), while proxy stimulation remained ineffective ($0.41\pm 0.38\leavevmode\nobreak\ \text{Hz}$, 80% reduction)
    • At $p_{\mathrm{inter}}=0.10$, closeness-centrality again prevailed ($7.69\pm 0.62\leavevmode\nobreak\ \text{Hz}$), but proxy stimulation showed relative improvement ($2.59\pm 1.48\leavevmode\nobreak\ \text{Hz}$, 66% reduction)
Figure 3: Strategic convergence in boosted networks. (a) Non-boosted networks exhibit distinct performance surfaces for top-centrality (solid color), proxy stimulation (shaded), and random (black) strategies. (b) Following centrality boosting, performance surfaces converge, indicating reduced efficacy differences between strategies. Shading denotes standard deviation across network ensembles.
Interpretive Synthesis The persistent >65% efficacy gap in proxy stimulation – even in boosted networks – underscores that spatial proximity to hubs cannot compensate for the absence of topological position. Optimal targets consistently exhibited high closeness centrality, reflecting their capacity to integrate information through short functional paths. By contrast, the modest improvement in proxy efficacy under high connectivity with boosting suggests that enhanced network integration partially compensates for suboptimal target selection.
Figure 4: Dynamic propagation characteristics. (A) Non-boosted networks: Top-centrality shows narrowband enhancement at driving frequency (10 Hz). (B) Boosted networks: All strategies exhibit spectral peaks but with varying signal-to-noise ratios (SNR: top=7.2 dB, hybrid=5.1 dB, proxy=2.8 dB).
4 Discussion
Our results demonstrate that topological centrality provides an effective criterion for driver node selection in modular neural networks. The superior performance of top-centrality nodes stems from their dual advantage: (1) high integration within the source population enables efficient input amplification, and (2) preferential connection to between-module bridges facilitates inter-population propagation.
The limited efficacy of proxy stimulation suggests that proximity to target hubs alone is insufficient for efficient control. This aligns with recent findings that effective control requires both local integration and global bridging capabilities [5]. Our spectral analysis further indicates that centrality-based selection preserves temporal fidelity of transmitted signals.
Notably, the performance gap between selection strategies was most pronounced at critical $p_{\mathrm{inter}}$ values (0.04-0.07), corresponding to the theoretical percolation threshold for modular networks. This suggests that topological control strategies are particularly valuable near phase transitions in network dynamics.
5 Conclusion
We establish that targeted stimulation of top-centrality neurons significantly outperforms both proximity-based and random selection strategies for activity propagation between neural populations. Our key findings demonstrate:
  • Critical operating regime: Centrality-based selection achieves maximum advantage ($\sim$2.7$\times$ efficiency gain) near network percolation thresholds ($p_{\mathrm{inter}}$ = 0.04-0.07), where biological neural systems typically operate
  • Metric-specific performance: Degree and closeness centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement
  • Mechanistic distinction: Spectral analysis revealed that centrality-based stimulation preserves temporal fidelity of transmitted signals, while proxy strategies amplify broadband noise
These results establish that effective neural control requires global topological influence rather than mere proximity to target hubs. The identified performance peak at critical connection densities provides a principled framework for precision neuromodulation in neurological disorders characterized by synchronization deficits. Future work should extend this approach to dynamic networks with adaptive topology and validate findings in in vivo recordings.
Acknowledgments
The study was supported by the Russian Science Foundation, project number 24-21-00470.